OpenStudy (anonymous):

Medal!! What is the fifth term of the geometric sequence? a1 = 120, a2 = 36, a3 = 10.8, a6 = 0.2916

OpenStudy (anonymous):

The sequence is, you simply multiply 3/10 to get a next term.

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

common ratio =a2/a1=36/120=3/10 \[a6=a5*\frac{ 3 }{10},a5=a6*\frac{ 10 }{ 3 }\]